# Cross-Problem Solving for Network Optimization: Is Problem-Aware Learning the Key?

### v1

@@ V-D2 Results
P1_D5 0.99 1.00 0.98
P1_D6 0.99 1.00 0.99
P1_D7 0.99 1.00 0.99
P2_D5 0.96 1.00 0.95
P2_D6 0.96 1.00 0.93
P2_D7 0.95 1.00 0.92
P3_D5 0.89 0.88 0.94
P3_D6 0.90 0.90 0.94
P3_D7 0.89 0.89 0.93
P4_D5 1.00 1.00 1.00
P4_D6 1.00 1.00 1.00
P4_D7 1.00 1.00 1.00
P5_D5 0.99 1.00 0.99
P5_D6 0.99 1.00 0.99
P5_D7 0.99 1.00 0.99
P6_D5 0.97 0.97 1.00
P6_D6 0.99 0.99 1.00
P6_D7 1.00 1.00 1.00
V-D3 Summary
In summary, the results presented in Fig. 7 and Fig. 8 demonstrate that incorporating constraint loss significantly improves the feasibility of the solutions generated by the PAD model, while also contributing to better overall solution quality. These findings validate the effectiveness of the constraint-aware module and highlight the potential of using constraint loss to enhance the feasibility of model outputs.
@@ VI Conclusion and Future Directions
This paper has demonstrated a problem-aware learning approach for cross-problem network optimization, aiming to provide general-purpose solution capabilities for fundamental models in wireless communication networks. Specifically, we have designed a problem-aware diffusion (PAD) model that explicitly incorporates the mathematical representations of problem formulations via problem embedding. This enables the model to recognize and understand the mathematical structure of optimization problems.
We have identified three key insights from experiments on six representative network optimization problems, which vary in type and complexity. First, the mathematical formulations of network optimization problems contain rich structural and semantic information. Even simple compressed representations can serve as effective and logically consistent problem feature embeddings. Second, explicitly introducing these embeddings during model training and inference significantly enhances cross-problem generalization, allowing the model to generate high-quality solutions even for previously unseen problems. Third, constructing an auxiliary constraint-aware module for identifying infeasible solutions and applying it as a loss function can effectively improve the feasibility of generated solutions, while also contributing positively to solution quality. These insights offer a foundational direction for building general-purpose AI models for wireless networks, especially in tasks such as network operation and resource management.
Additionally, several aspects of the current work warrant further investigation. From an engineering perspective, multi-task learning, model compression, and real-world deployment of the PAD model remain essential considerations. On a deeper level, the current PAD framework does not yet incorporate mathematical analysis of different optimization problem characteristics, nor does it learn corresponding mathematical reasoning strategies. We believe that new neural architecture designs and optimization-specific learning objectives are necessary to further enhance the effectiveness of problem-aware learning.

### v2

@@ V-D3 Summary
Model Parameter Quantity 2.04 Million 52 Trillion
Runtime Processor Utilization 48% (GPU) $10\%\sim 95\%$ (CPU)
Runtime Memory Usage 1.55GB (GPU) 0.3GB (CPU)
[…]
TABLE IV: The impact of removing one problem from the four training problems on the PAD model’s GLOB_GT_GAP $\downarrow$ performance. Each column corresponds to the GLOB_GT_GAP performance on a different cluster or an individual problem. For models trained with one problem removed, a (+) indicates that the score is better than that of the full training model, while a (–) indicates it is worse.
P1, P2, P3 P4, P5, P6 P7, P8, P9 P10
Trained on P1, P4, P7, P10 $7.46\times 10^{-2}$ $7.17\times 10^{-2}$ $10.63\times 10^{-2}$ $20.04\times 10^{-2}$
Trained on P4, P7, P10 $7.52\times 10^{-2}$ (-) $7.62\times 10^{-2}$ (-) $5.37\times 10^{-2}$ (+) $21.20\times 10^{-2}$ (-)
Trained on P1, P7, P10 $8.85\times 10^{-2}$ (-) $7.63\times 10^{-2}$ (-) $6.94\times 10^{-2}$ (+) $32.95\times 10^{-2}$ (-)
Trained on P1, P4, P10 $8.94\times 10^{-2}$ (-) $7.14\times 10^{-2}$ (+) $12.90\times 10^{-2}$ (-) $28.87\times 10^{-2}$ (-)
Trained on P1, P4, P7 $8.86\times 10^{-2}$ (-) $5.78\times 10^{-2}$ (+) $10.06\times 10^{-2}$ (+) $24.10\times 10^{-2}$ (-)
V-E PAD Efficiency Analysis
V-E1 Specific Settings
To demonstrate the superior cross-problem generalization efficiency of PAD’s single-model approach compared with other methods, we introduce the LLM-based solver OPRO [38] and the traditional numerical solver Gurobi [59], and conduct comparisons in terms of solution quality, algorithm construction cost, and runtime cost. Given that many existing works [60, 6, 33] have already established the advantages of GDM-based solution generators over other discriminative models and DRL methods, and that PAD is specifically designed to explore cross-problem generalization, we do not include such baselines in this section. Furthermore, we extend Table I by increasing the number of parallel samples in PAD, in order to showcase its solution quality and efficiency under different $p_{num}$ settings.
V-E2 Results
As shown in Table III, while PAD does not match the optimal solver Gurobi in terms of solution quality and feasibility, it outperforms OPRO in most cases. This may partly stem from some problems lying beyond the knowledge scope of GPT-5. In terms of training or running efficiency, PAD is far more efficient and controllable in terms of hardware, computational resources, and time consumption. OPRO requires a massive number of model parameters and performs poorly on certain problems. Since GPT-5 is accessed via an API, its runtime hardware metrics are unavailable; however, its computational cost is expected to exceed that of PAD. Gurobi, on the other hand, relies heavily on manual coding and debugging, with solver construction typically requiring several hours per problem. Moreover, Gurobi is entirely incapable of handling problems outside its predefined design scope. In terms of inference efficiency, PAD substantially surpasses OPRO, whose model size and computational demands are disproportionate to the problem complexity, and it also consistently achieves lower latency than OPRO. Compared with Gurobi, PAD delivers comparable runtime performance while operating on common hardware platforms.
As shown in Fig. 9, we evaluate the impact of parallel sampling on solution quality under four configurations: $p_{num}=2,4,8,16$. Increasing $p_{num}$ significantly improves the final solution quality, approaching optimal performance on many problems at $p_{num}=16$, as this enables exploration of a broader distribution of high-quality solutions. In terms of efficiency, the four parallel sampling settings achieve average inference times of 117.25ms, 183.98ms, 315.31ms, and 600.47ms for single solution generation, respectively. These runtimes do not scale linearly with $p_{num}$ because the model’s nonlinear matrix operations prevent simple doubling of runtime with doubled sampling.
V-E3 Summary
Overall, PAD demonstrates superior efficiency in model construction cost, inference cost, and inference latency compared to existing LLM-based solvers. Moreover, while its performance is already comparable to traditional numerical solvers in many aspects, PAD offers a more efficient and scalable potential advantage, particularly in cross-problem generalization.
V-F Analysis of Training Problem Selection
V-F1 Specific Settings
To analyze how the choice of training problems affects the model’s cross-problem generalization, we conduct four ablation studies by removing one problem at a time from the complete set of four training problems. In addition, this setup also serves an engineering purpose: to explore how clustering problems according to common mathematical properties influences the model’s within-cluster generalization performance.
V-F2 Results
As shown in Table LABEL:tab_training_problem_ablation, we use the symbols (+) and (-) to indicate whether the performance of a given model is better or worse than that of the baseline model trained on the complete set of problems (the first row). From the diagonals of the last four rows, it is clear that the within-cluster generalization performance consistently degrades when the problem from that cluster is removed. This suggests that the excluded training problem indeed provides transferable knowledge for cross-problem generalization within its cluster, thereby supporting the validity of clustering problems according to common mathematical properties. Fundamentally, this also indicates that network optimization problems grouped by such mathematical properties share overlapping or transferable structures in the distribution of high-quality solutions.
For the clusters corresponding to the problems that were not removed, we argue that the observed performance gains or drops primarily result from the interplay between shared knowledge and competition for limited model capacity. In Table LABEL:tab_training_problem_ablation, the first and fourth columns both show degraded performance across all ablation models, indicating that problems in these clusters share substantial knowledge with those in other clusters (reflected at the neural network level by complementary gradient directions across problem). Removing training problems in these clusters reduces the availability of such shared knowledge, and the additional free model capacity is insufficient to compensate for the loss. Intuitively, both the first and fourth columns correspond to convex-like optimization problems, whose relatively simple high-quality solution distributions are less sensitive to model capacity than to shared knowledge. The second and third columns follow a similar pattern, but their performance is more strongly constrained by model capacity rather than shared knowledge. This is likely because the high-quality solution distributions of non-convex optimization, mixed-integer programming, and combinatorial optimization problems are inherently more diverse and complex.
V-F3 Summary
The choice of training problems plays a critical role in determining the cross-problem generalization performance of the PAD model. Selecting representative problems from the same category based on common mathematical classifications can enhance the model’s ability to generalize to other problems within that category, while more complex problems also impose higher demands on model capacity. Moreover, problems grouped into different categories may still exhibit varying degrees of shared knowledge.
@@ VI Conclusion and Future Directions
This paper has demonstrated a problem-aware learning approach for cross-problem network optimization, aiming to provide a promising new direction for equipping future foundation models in wireless networks with cross-problem solving capabilities. Specifically, PAD explicitly incorporates the mathematical representations of problem formulations via problem embedding, enabling the model to capture the mathematical structure of the optimization problems.
We have identified three key insights from experiments on ten representative network optimization problems, which vary in type and complexity. First, the mathematical expression–based feature embedding of network optimization problems contains rich structural and semantic information. When used as explicit conditional input, it enables a single model to achieve effective cross-problem generalization and generate high-quality solutions even on problems unseen during training. Second, constructing an auxiliary constraint-aware module to identify infeasible solutions and incorporating it as a loss function can not only improve the feasibility of generated solutions but also enhance solution quality. Third, clustering problems by their mathematical characteristics and selecting only representative problems from each cluster for training supports cross-problem generalization within each cluster. Furthermore, partial knowledge sharing across different clusters can also contribute to performance gains. Notably, with these three aspects realized, PAD surpasses LLM-based solvers in efficiency (such as algorithm construction and inference latency) while matching advanced numerical solvers, and simultaneously offers greater flexibility and scalability potential.
Additionally, several aspects of the current work warrant further investigation. First, the relationship between problem-feature embedding and solution generation. Although the current embedding strategy achieves a certain degree of cross-problem generalization, compared with tasks such as text-to-image generation or trajectory control, the linkage between training loss and solution quality through problem feature embedding remains relatively weak. Strengthening the model’s semantic understanding of problem feature embedding, as well as its direct influence on solution quality, is therefore essential. Second, improving the feasibility of generated solutions more efficiently remains a challenge. The current reliance on an external constraint-aware module and constraint loss still leaves room for improvement in terms of both efficiency and determinism. Third, optimizing model architecture and size is also necessary. Given the observed competition among multiple problems for model capacity, more compact yet effective neural designs are needed. Fourth, the classification and boundary of problems deserve further exploration. A deeper analysis of the transferability of high-quality solution distributions across objective functions is required to better clarify the limits of cross-problem generalization and to clarify how those boundaries may be expanded.